26. A company's income before interest expense and income taxes is $100,000 and its interest expense is $50,000. Its times interest earned ratio is:

  • 0.50

  • 1.50

  • 2.00

  • 1.00

  • 0.70

    27. A registrar keeps stockholder records and prepares official lists of stockholders and dividend payments.

  • true of false

    29.Eastline Corporation had 16,500 shares of $5 par value common stock outstanding when the board of directors declared a stock dividend of 5,445 shares. At the time of the stock dividend, the market value per share was $15. The entry to record this dividend is:

  • No entry is needed.

  • Debit Retained Earnings $27,225; credit Common Stock Dividend Distributable $27,225.

  • Debit Common Stock Dividend Distributable $81,675; credit Retained Earnings $81,675.

  • Debit Retained Earnings $81,675; credit Common Stock Dividend Distributable $81,675.

  • Debit Retained Earnings $81,675; credit Common Stock Dividend Distributable $27,225; credit Paid-In Capital in Excess of Par Value, Common Stock $54,450.

    31. Peavey Enterprises purchased a depreciable asset for $22,000 on April 1, Year 1. The asset will be depreciated using the straight-line method over its four-year useful life. Assuming the asset's salvage value is $2,000, Peavey Enterprises should recognize depreciation expense in Year 2 in the amount of:

  • $10,000

  • $5,000

  • $5,500

  • $20,000

  • $9,250

  • 39 -The amount due on the maturity date of a $11,200, 90-day 7%, note receivable is: (Use 360 days a year.)

  • $11,396.

  • $11,200.

  • $10,416.

  • $11,984.

  • $11,004.

    40-A company had average total assets of $937,000. Its gross sales were $1,099,000 and its net sales were $960,000. The company's total asset turnover equals:

  • 1.14.

  • 1.02.

  • 0.98.

  • 0.85.

  • 1.17.
Answer #1

26)correct option is "C" -2

Times interest earned ratio = Income before interest and tax /Interest expense

                             = 100000/ 50000

                             = 2

27)The statement is "True.

Registrar keeps stockholders records and prepares official lists of stockholders.

29)correct option is"B"

% of stock dividend =shares issued as stock dividend /shares outstanding

                    = 5445/16500

                   = .33 or 33%

So this is a case of large stock dividend .The journal entry includes

Debit Retained earning and credit common stock distributable with (par value *stock dividend issued) = 5*5445= 27225

31)correct option is "B"

Depreciation expense for year 2 =[Cost- salvage value ]/useful life

                        =[22000-2000]/4

                        = 20000/4

                        = $5 000